OpenStudy (anonymous):

Find the derivative of f(x) = -2x2 + 11x at x = 9

OpenStudy (bibby):

\(\dfrac{d}{dx}-2x2 + 11x=\dfrac{d}{dx}-2x2 + \dfrac{d}{dx}11x=\)

OpenStudy (anonymous):

@bibby I do not understand

OpenStudy (bibby):

take the derivative of each term and put them together the derivative of f(x) = -2x2 + 11x is equal to the derivative of -2x2 plus the derivative of 11x